OpenStudy (anonymous):

how can i find the area under the graph ln(x) using integral function? can somebody help me, and thanks in advance.

OpenStudy (jack1):

sorry, what is the function of the line? ie f(x) = ln(x) ? and between what x coordinates / limits do u want the area to be? ie max x value, min x value...?

OpenStudy (jack1):

...? @iraaaaa

OpenStudy (anonymous):

sorry for the vague question. for example, ln(x)=∫max value is 2, and min value is 1. but my equation involves exponential

OpenStudy (jack1):

ok, so write the exponential equation here and we'll work it out together

OpenStudy (anonymous):

\[\ln(x)= \int\limits_{1}^{2} (1/t) dt\]

OpenStudy (jack1):

ok, so integral of 1/t = ln(t) + a constant constant is negated in the presence of limits

OpenStudy (jack1):

integral of 1/t = ln(t) + constant \[\int\limits_{1}^{2} (1/t) = [\ln(t) + C]\lim_{2 \rightarrow 1}\] \[= [\ln(2) +C] - [\ln(1) +C] = \ln2 - \ln1\]

OpenStudy (jack1):

= 0.6932...ish = ln(x)

OpenStudy (jack1):

i'm not sure if i'm helping here... @ganeshie8 u got a sec?

OpenStudy (anonymous):

do we have to insert the constant C and the limit even if they're probably not needed?

OpenStudy (jack1):

no, its just good to get in the practice to always remember to write the canstant in as you're not always dealing with limits

OpenStudy (jack1):

could you please look over this @.Sam. Im not sure?

OpenStudy (jack1):

by integrating, you are working out the area under the graph

OpenStudy (anonymous):

so, integrating will give out the area? i'm gettin it. btw, what's the difference between 'ln' and 'e' ?

OpenStudy (jack1):

ln is natural log = log to base e e = 2.71828182845904...etc, its a mathematical constant

OpenStudy (jack1):

they're kinda like opposing opporators, think plus and minus + opposite of - so if u had e^10 the natural log of e^10 = ln e^10 = 10

OpenStudy (jack1):